OpenStudy (anonymous):

Use linear approximation, i.e. the tangent line, to approximate the square root of 49.2 as follows: Let f(x)=x. The equation of the tangent line to f(x) at x=49 can be written in the form y=mx+b where m = ? b = ?

OpenStudy (anonymous):

49 + (x-49) = x m = 1 b = 0

OpenStudy (anonymous):

I'm sorry that was my bad. The question actually is: Use linear approximation, i.e. the tangent line, to approximate the square root of 49.2 as follows: Let f(x)= the square root of x. The equation of the tangent line to f(x) at x=49 can be written in the form y=mx+b where m = ? b = ?

OpenStudy (anonymous):

f(a) + f'(a) (x-a) In this case, a = 49

OpenStudy (anonymous):

Ok... So then what do I do?

OpenStudy (anonymous):

what is the derivative of square root (x) ?

OpenStudy (anonymous):

1 / (2 * square root(x))

OpenStudy (anonymous):

square root (7) = plus / minus 7

OpenStudy (anonymous):

\[\pm 7 + \frac{1}{\pm 14} (x-49)\]

OpenStudy (anonymous):

Okay so what am I supposed to do next?

OpenStudy (anonymous):

\[m = \frac{1}{\pm 14}\]

OpenStudy (anonymous):

Got it :) There's one more part to the problem, it says: Using this (the answers from the previous parts), we find our approximation for square root (49.2) = ?

OpenStudy (anonymous):

It also says: NOTE: For this part, give your answer to at least 9 significant figures or use fractions to give the exact answer.

OpenStudy (anonymous):

plug in 49.2 for x